OpenStudy (anonymous):

Which statement is true about the equation fraction 3 over 4z = fraction 1 over 4z − 3 + 5?

OpenStudy (texaschic101):

3/4z = 1/4z - 3 + 5 3/4z = 1/4z + 2 what are your answer choices ??

OpenStudy (texaschic101):

3/4z = 1/4z + 2 3/4z - 1/4z = 2 2/4z = 2 z = 2/(2/4) z = 2 * 4/2 z = 8/2 which reduces to 4 well....z = 4...thats true